Shift report 20000705 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Jul 05 2000 - 00:43:00 EDT

  • Next message: Flemming Videbaek: "Fw: July 10 RHIC lunchtime discussion"

    ------------------------------------------------------------
    Date: Wed Jul  5 00:43:00 2000, Shift: 16:00-24:00
    Supervisor: ko, with: ma, rami, zis
    Log book#2 pages 134 - 135
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine was sick all shift, going through the cycle:
       - inject (intensity as high as 2500+2500 in both beams)
       - start ramp
       - immediately loose one beam
       - loose the second beam at transition
       - ramp down,
       - repeat every 30-40 minutes.
    
    2. Fire alarm in our area (1002B). Had a visitation
       by firemen (2 fire trucks + 3 aux cars), BNL police,
       HP and MCR. Firemen went into tunnel towards the 1003
       area, did not find any fire, alarm cleared itself.
       I do not think firemen (with big axes, etc) went
       into the 1002 IR. Conclusion: false alarm.
    
       Current machine status: resweeping tunnel (sweep lost
       by firemen), also yellow quench link interlock.
    
       Further beam highly unlikely, experiment was shutdown
       (TPC high-voltages are off,
       BB, Tiles, H1/H2, etc... hv left on,
       magnets are off), everybody goes home and there
       is no 24-to-08 shift.
    
    3. No known detector problems.
    
    K.O.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Jul 05 2000 - 00:46:54 EDT